A Plane flies from Portland, ME to El Paso, TX with an average speed of
303mph. A car drives the same distance and leaves at the same time, averaging
a speed of 50 mph. The plane arrives in El Paso 35.42 hours before the car.
How far is it from Portland to El Paso?
Please Help
E

Found 2 solutions by asha, ptaylor:
Answer by asha(30)

let the distance from Portland to El Paso be x miles
the speed of the plane is 303 m.p.h
the speed of the car is 50 m.p.h.
time taken by the plane to fly the distance is x/303 hrs.
time taken bythe car to travel the same distance is x/50 hrs.
but it is given that
x/50 -x/303 =35.42
x(1/50-1/303) =35.42
x(303-50)303*50=35.42
253*x =35.42*303*50
x=( 35.42*303*50)/253
= 2121 miles
you can check the answer by plugging in the values in the equation.
good luck !!!

Answer by ptaylor(1894)

Let x=Distance from Portland to El Paso
Now the time it takes the plane to make the trip is the distance (x) divided by the rate (303 mph) and , likewise, the time it takes the car to make the trip is distance divided by rate which is (x) divided by 50 mph.
We know that the difference between these two times is 35.42 hours. Thus, the equation to solve is:
x/50-x/303=35.42 Lets clear the fractions by multiplying both sides of the equation by(50)(303). This gives us:
303x-50x=(35.42)(50)(303) simplifying, we get:
253x=536,613
x=2121 miles

Hope this helps------ptaylor